Tutor could you help me with this problem please?
Find a linear functon g given g(-1/4)= -6 and g(2)=3
Thank you

Answer by stanbon(75887)

Find a linear functon g given g(-1/4)= -6 and g(2)=3
--------------------
You have two Points: (-1/4,-6) and (2,3)
----------------------
slope = (3--6)/(2--1/4) = 9/(9/4) = 4
----------------------------------------
Form is y = mx + b
You know y,x and m ; solve for "b":
3 = 4*2 + b
b = -5
---------------
EQUATION:
y = 4x -5
===============
Cheers,
Stan H.